# Positive intermediate Ricci curvature on cohomogeneity one manifolds in low dimensions

### v1

@@ 4.1. Dimension 5
$\{(1,1)\}\subset\{(e^{x\theta},1)\},\{(e^{yd\theta},e^{i\theta})\}\subset S^{3}\times S^{1},\vspace{0.1cm}$
@@
$\{(1,1)\}\subset\{e^{xp\theta},e^{i\theta}\},\{e^{yp\theta},e^{i\theta}\}\subset S^{3}\times S^{1},\vspace{0.1cm}$
where $x,y\in\mathrm{Im}(\mathbb{H})\cap S^{3}$. Moreover, the generators of the Weyl group are given by $w_{-}=w_{+}=(1,-1)$ or $(-1,-1)$, depending on whether $p$ is even or odd, and thus the geodesic $\gamma$ has length $2L$. Therefore, the space $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{}{\text{for some}}~{}t\}$ is two-dimensional if $x\neq y$ and one-dimensional otherwise. ∎
In the next example, we show that Proposition 4.3 does not generalize to the case in which the principal isotropy group has order two.
@@
We compare Proposition 4.3 to the metric with $\operatorname{Ric}_{2}>0$ on $S^{3}\times S^{2}$ described in Example 2.4. Identify $S^{2}$ with $S^{3}/S^{1}$, and consider the cohomogeneity one action of $S^{3}\times S^{1}$ on $S^{3}\times S^{2}$ given by
@@ 4.3. Dimension 7
In this final section, we first prove the obstruction statement in Theorem 1.1, then we complete the proofs of the remaining cases in Theorem 1.3.
@@
Denoting $H=\langle(i,i)\rangle$, we have that $N(H)=\{(e^{i\theta},e^{i\theta})\}\cdot\langle(j,j)\rangle$, and hence $N(H)_{0}=\{(e^{i\theta},e^{i\theta})\}$. Letting $K_{-}=\{(e^{ip_{-}\theta},e^{iq_{-}\theta})\}$ and $K_{+}=\{(e^{ip_{+}\theta},e^{iq_{+}\theta})\}\cdot H$, for any $n\in N(H)_{0}$, we have $nK_{-}n^{-1}=K_{-}$ and $nK_{+}n^{-1}=\{(e^{\ell p_{+}\theta},e^{\ell q_{+}\theta})\}\cdot H$ for some $\ell=(\cos\phi)j+(\sin\phi)k$. Hence any group diagram equivalent to $H\subset K_{-},K_{+}\subset G$ is of the form (up to switching $K_{-}$ and $K_{+}$):
@@
$w_{+}=\begin{cases}(-1,1)\cdot H^{\prime}&\text{if }p_{+}\not\equiv q_{+}\mod 2,\\ (v,y)\cdot H^{\prime}&\text{if }p_{+},q_{+}\text{ are odd and }p_{+}\equiv q_{+}\mod 4,\\ (-v,y)\cdot H^{\prime}&\text{if }p_{+},q_{+}\text{ are odd and }p_{+}\not\equiv q_{+}\mod 4.\end{cases}$
@@
$G_{\gamma(L)}$ =$=\{(e^{vp_{+}\theta},e^{yq_{+}\theta})\}\cdot H^{\prime},$
$G_{\gamma(3L)}$ =$=\{(e^{wp_{+}\theta},e^{zq_{+}\theta})\}\cdot H^{\prime}.$
@@
where $x,y\in\mathrm{Im}(\mathbb{H})\cap S^{3}$. For the case in which $n=1$, one has $w_{+}=(-1,1)$. Moreover, depending on the parity of $p$, $q$, $p/{\gcd(p,q)}$, and $q/{\gcd(p,q)}$, $w_{-}$ is one of $(1,-1)$, $(-1,1)$, or $(-1,-1)$. In all of the cases, the geodesic $\gamma$ has length equal to either $2L$ or $4L$. Furthermore, there are two distinct singular isotropy groups along $\gamma$, one of which has dimension one and the other one has dimension three. Thus the space $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{}{\text{for some}}~{}t\}$ is four-dimensional. This completes the proof in the case where $n=1$.
@@
$\{(\alpha e^{i\phi}\alpha^{-1},1,1)\}\subset\{(\alpha e^{i\phi}\alpha^{-1},\beta e^{ib\theta}\beta^{-1},e^{i\theta})\},S^{3}\times 1\times 1\subset S^{3}\times S^{3}\times S^{1}\vspace{0.02cm}$
for some $\alpha,\beta\in S^{3}$. For such a group diagram, one has $w_{-}=(1,\pm 1,-1)\cdot H$ and $w_{+}=(\alpha j\alpha^{-1},1,1)\cdot H$. Therefore, the geodesic $\gamma$ has length $4L$ and every singular isotropy group along $\gamma$ equals either $G_{\gamma(0)}$ or $G_{\gamma(L)}$. Thus the space $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{}{\text{for some}}~{}t\}$ is three-dimensional and the claim follows. ∎

### v2

@@ 4.1. Dimension 5
$\{(1,1)\}\subset\{(e^{x\theta},1)\},\{(e^{yd\theta},e^{i\theta})\}\subset S^{3}\times S^{1},\vskip 2.84544pt$
@@
$\{(1,1)\}\subset\{e^{xp\theta},e^{i\theta}\},\{e^{yp\theta},e^{i\theta}\}\subset S^{3}\times S^{1},\vskip 2.84544pt$
where $x,y\in\mathrm{Im}(\mathbb{H})\cap S^{3}$. Moreover, the generators of the Weyl group are given by $w_{-}=w_{+}=(1,-1)$ or (1,1)(-1,-1), depending on whether $p$ is even or odd, and thus the geodesic $\gamma$ has length $2L$. Therefore, the space $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{\text{for some}}~t\}$ is two-dimensional if $x\neq y$ and one-dimensional otherwise. ∎
The manifolds in the family $Q_{A}^{5}$ are diffeomorphic to $S^{3}\times S^{2}$; see [Hoe10, Table VII]. In the next example, we show that $S^{3}\times S^{2}$ with the action described in Example 2.4 is a member of $Q_{A}^{5}$, and the fact that it admits an invariant metric with $\operatorname{Ric}_{2}>0$ shows that Proposition 4.3 does not generalize to the case in which the principal isotropy group has order two.
@@
Identify $S^{2}$ with $S^{3}/S^{1}$, and consider the cohomogeneity one action of $S^{3}\times S^{1}$ on $S^{3}\times S^{2}$ given by
@@ 4.3. Dimension 7
In this final section, we first prove the obstruction statement in Theorem A, then we complete the proofs of the remaining cases in Theorem C.
@@
Denoting $H=\langle(i,i)\rangle$, we have that $N(H)=\{(e^{i\theta},e^{i\theta})\}\cdot\langle(j,j)\rangle$, and hence $N(H)_{0}=\{(e^{i\theta},e^{i\theta})\}$. Letting $K_{-}=\{(e^{ip_{-}\theta},e^{iq_{-}\theta})\}$ and $K_{+}=\{(e^{jp_{+}\theta},e^{jq_{+}\theta})\}\cdot H$, for any $n\in N(H)_{0}$, we have $nK_{-}n^{-1}=K_{-}$ and $nK_{+}n^{-1}=\{(e^{\ell p_{+}\theta},e^{\ell q_{+}\theta})\}\cdot H$ for some $\ell=(\cos\phi)j+(\sin\phi)k$. Hence any group diagram equivalent to $H\subset K_{-},K_{+}\subset G$ is of the form (up to switching $K_{-}$ and $K_{+}$):
@@
[…]
@@
$G_{\gamma(L)}$ $=\{(e^{vp_{+}\theta},e^{yq_{+}\theta})\}\cdot H^{\prime},$
$G_{\gamma(3L)}$ $=\{(e^{wp_{+}\theta},e^{zq_{+}\theta})\}\cdot H^{\prime}.$
@@
where $x,y\in\mathrm{Im}(\mathbb{H})\cap S^{3}$. For the case in which $n=1$, one has $w_{+}=(-1,1)$. Moreover, depending on the parity of $p$, $q$, $p/{\gcd(p,q)}$, and $q/{\gcd(p,q)}$, $w_{-}$ is one of (1,1)(1,-1), (1,1)(-1,1), or (1,1)(-1,-1). In all of the cases, the geodesic $\gamma$ has length equal to either $2L$ or $4L$. Furthermore, there are two distinct singular isotropy groups along $\gamma$, one of which has dimension one and the other one has dimension three. Thus the space $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{\text{for some}}~t\}$ is four-dimensional. This completes the proof in the case where $n=1$.
@@
$\{(e^{x\phi},1,1)\}\subset\{(e^{x\phi},e^{yb\theta},e^{i\theta})\},S^{3}\times 1\times 1\subset S^{3}\times S^{3}\times S^{1}\vskip 0.56917pt$
for some $x,y\in\mathrm{Im}(\mathbb{H})\cap S^{3}$. For such a group diagram, one has $w_{-}=(1,\pm 1,-1)\cdot H$ and $w_{+}=(z,1,1)\cdot H$ for some $z\in\mathrm{Im}(\mathbb{H})\cap S^{3}$ such that $z$ is orthogonal to $x$. Therefore, the geodesic $\gamma$ has length $4L$ and every singular isotropy group along $\gamma$ equals either $G_{\gamma(0)}$ or $G_{\gamma(L)}$. Thus $\operatorname{span}\{J\in\Lambda\mid J(t)=0~{\text{for some}}~t\}$ is three-dimensional and the claim follows. ∎